One alloy contains 80% gold (by weight)and another alloy contains 55% gold.how many grams of each alloy should be combined to make 40g of an alloy that contains 70% gold.
Answer by Alan3354(69443)

One alloy contains 80% gold (by weight)and another alloy contains 55% gold.how many grams of each alloy should be combined to make 40g of an alloy that contains 70% gold.
===================================
e = amount of 80%
f = amount of 55%
------
e + f = 40 (total alloy)
80e + 55f = 70*40 (total gold)
------------